### v1

# Two-Stage Stance Labeling: User-Hashtag Heuristics with Graph Neural Networks thanks: Army Research Office

Abstract
The high volume and rapid evolution of content on social media present major challenges for studying the stance of social media users. In this work, we develop a two stage stance labeling method that utilizes the user-hashtag bipartite graph and the user-user interaction graph. In the first stage, a simple and efficient heuristic for stance labeling uses the user-hashtag bipartite graph to iteratively update the stance association of user and hashtag nodes via a label propagation mechanism. This set of soft labels is then integrated with the user-user interaction graph to train a graph neural network (GNN) model using semi-supervised learning. We evaluate this method on two large-scale datasets containing tweets related to climate change from June 2021 to June 2022 and gun control from January 2022 to January 2023. Experiments demonstrate that our user-hashtag heuristic and the semi-supervised GNN method outperform zero-shot stance labeling using LLMs such as GPT4. Further analysis illustrates how the stance labeling information and interaction graph can be used for evaluating the polarization of social media interactions on divisive issues such as climate change and gun control.
Index Terms:
graph neural networks, polarization, social network analysis, stance labeling
I Introduction
In today’s digital society, online social media have become the main platform for dissemination of partisan and political information as well as the main forum for public discussion of political and social issues. The immense scale and interconnected nature of social media has meant that the scope of public discussion online permits varied and diverse viewpoints about a broad range of topics. Despite the potential for the co-existence of diverse perspectives online, research has shown that online communities are increasingly polarized, especially with respect to contentious issues [1, 2, 3, 4]. The increasingly polarized nature of online communities poses a real challenge to many democracies in today’s world. Studies have shown that over the last 30 years, both Democrats and Republicans have become more negative in their views towards the opposition party [4], and these negative views have affected outcomes in areas such as scholarship fund allocation, mate selection, employment decisions, and acceptance of public health measures [5, 6]. Over the last several years, social media platforms like Twitter, Reddit, YouTube, and others have amassed millions of users who engage both with mainstream and fringe media outlets as well as engaging directly with other users on numerous political and social issues. As such, large social media platforms like Twitter present an excellent opportunity to study web-scale user behavior and polarization among differing viewpoints on contentious issues such as climate change and gun control.
The study of polarization online has broadly been divided into two main categories. The majority of computational social science research has focused on analyzing interactional polarization, which occurs when users interact to a large extent exclusively with like-minded individuals in highly balkanized communities and are only minimally exposed to users with opposing viewpoint. Such echo chambers isolate users in an in-group community that reinforces users’ existing beliefs and skews user perspectives towards a particular issue [7]. Another important form of polarization that is less well-studied is known as affective polarization. Affective polarization refers to highly negative sentiments expressed by users towards out-groups of users with opposing viewpoints [3, 1]. Prior work on affective polarization has been limited in part due to reliance on manually annotated datasets and challenges in identifying user stances and group memberships in a computationally efficient manner. With recent advances in language modeling, stance detection, sentiment analysis, and network science-based approaches to community detection, researchers have begun to study affective polarization using large, unannotated corpora and interaction networks. Despite this recent progress, much work on affective polarization remains ad hoc and incommensurate between studies.
Stance labeling remains a particularly challenging task due to the informal and noisy nature of social media language, the presence of sarcasm and irony, and the potential biases introduced by human annotators and automated models. To further facilitate research into polarization online, we develop a stance labeling method that incorporates both textual content and user social interactions in a scalable two-stage pipeline. In the first stage, we apply a reciprocal label propagation algorithm to the user-hashtag bipartite graph that generates a soft-labeled set of users. In the second stage, we construct a signed, weighted, and attributed user-user interaction network using transformer language models to embed the content of user posts and to determine the sentiment of interactions between users. Using the soft labels generated from the label propagation stage, we employ semi-supervised training to train a graph neural network classifier to classify user stances.
Our stance labeling approach utilizes an understanding about linguistic and network homophily. Linguistic homophily states that people who share similar ideologies also share similar textual content, not only through similar keyword and hashtag usages but also through sentiment and linguistic features [8, 9], and the inverse phenomenon-that users with dissimilar views use different language-has been investigated by researchers [2]. Likewise, the concept of network homophily, which describes the phenomenon that similar nodes tend to be connected to one another has long been studied in graph theory [10, 11] and informs the success of graph neural networks [12, 13, 14]. By combining the content-based and network-based approaches into a single, scalable pipeline for stance labeling, we facilitate further large-scale analysis of both interactional and affective polarization online.
In sum, the contributions of this work are:
  • We develop an approach for estimating user stance labels that incorporates both textual features and social network interactions
  • A two-stage pipeline allows our approach to scale to large datasets with soft labeling and semi-supervised training
  • We evaluate our method on two large-scale datasets on two contentious political issues: climate change and gun control
II Related Work
As social media has become the primary forum for discussion on many topics, including political issues, there has been growing interest in user stance detection . Many existing methods can be divided into two broad categories: textual-based methods that use the content of messages, and network-based methods that use social or conversational structure. Textual-based methods use the content of tweets, user profiles, and hashtags to determine a user’s stance on a particular topic. Several studies have used supervised classification of user stance labels. Classifiers are trained using a set of features encoding tweet text, hashtags, and user profile information. These techniques have been applied to classify user political ideology . These approaches are generally rely on manually annotated sets of users to train the classifier models, and such methods are often limited by the size and diversity of the training content as manual annotation is both time-consuming and expensive. Furthermore, these methods can be susceptible to annotator bias poor generalizability due to the limited distribution of training examples.
Topic modeling and user clustering are two unsupervised techniques that have been employed for stance labeling. These unsupervised textual generally embed tweets or user profile information in a latent feature space and then apply a variety of topic modeling or cluster methods to identify tweets of similar topics or clusters of similar users. While such techniques avoid the issues with manual annotation for supervised training, they suffer from the  curse of dimensionality—the search space for a solution grows exponentially with the increase in dimensionality as there are many more possible patterns than in lower-dimensional subspaces. Selecting useful and informative features, and the computation and memory complexity of clustering methods on high-dimensional create problems for unsupervised clustering-based methods. For high-dimensional data, many clustering techniques fail to produce meaningful clusters, but on the other hand, such techniques can be very efficient on low-dimensional features spaces . Several studies have therefore employed dimensionality reduction techniques prior to clustering to first project tweets to a low-dimensional space, and such techniques can be robust to class imbalance between different stance groups.
Network-based approaches use the structure of social interaction networks to determine user ideology or stance. Such methods leverage techniques such as label propagation and min-cut, and node embedding techniques to leverage structural information from retweet  and user follower  networks in order to train. Xiao et al.  generated a multi-relational network using retweets, mentions, likes, and follows for binary classification of user ideologies.
More recent works have combined both textual content and network features in socially-infused text mining. Li and Goldwasser  combine user interactions and user sharing of news media to detect bias in news articles, and Johnson, Jin, and Goldwasser  combine multiple features from lexical features and social interactions to categorize tweets by politicians on a varity of political topics. Pan et al.  use the network structure and node content to learn node representations for classifying the category of scientific publications. Such works have demonstrated improvements over earlier methods that rely solely on textual content or network structure alone.
Our approach is similar in its reliance on socially-infused text mining for stance labeling, but our approach is unique from the methods described above in that we i) combine language features from state-of-the-art transformer language models with the structure and sentiment of social interactions using graph neural networks, and ii) our two-stage process allows our method to scale to large datasets with minimal supervision.
III Methods
This section describes our proposed method to estimate the stance of users as a binary classification problem. We first describe the collection of data for the gun control and climate changes datasets. Then we characterize the hashtag-based heuristic method to generate soft labels for two stance groups of users, which are used as seed users for training and evaluating stance labeling methods. We then introduce several baseline methods followed by our GNN-based approach for stance labeling.
III-A Data Collection
We collect tweets related to two prominent and contentious political issues—climate change and gun control. The keywords listed in Table  I are used to scrape the initial set of relevant tweets related to each topic. To fully capture conversation cascades and user interactions, we then recursively collect all referenced tweets from each tweet in the initially matched set. This ensures that all available tweets from relevant conversations are included in our analysis, even if particular tweets do not explicitly use any of the keywords. The climate change dataset consists of tweets published between June 1st, 2021 and May 31st, 2022. In all, the climate change dataset consists of a total of 46M tweets authored by 4.8M unique users and contains conversation threads of at least three tweets. The gun control dataset consists of tweets published between January 1st, 2022 and December 31st, 2022. In all, the gun control dataset dataset consists of a total of 14.4M tweets from 2.66M unique users containing 335,000 conversation threads of at least three tweets.
TABLE I: Set of keywords used to collect tweets related to Climate Change and Gun Control from Twitter.
Climate Change Gun Control
climate change gun control
global warming gun rights
climate hoax second amendment
global warming hoax 2nd amendment
global cooling #guncontrol
#ActOnClimate #guncontrolnow
#ClimateChange #gunreform
#climatechangehoax #gunviolence
#globalwarminghoax #endgunviolence
#globalcooling #2a
#globalwarmingisahoax #nra
#climatehoax #gunrights
#secondamendement
#shallnotbeinfringed
#righttobeararms
Not only are these two topics highly relevant political issues that generate a large amount of engagement, news coverage, and user interaction on social media, but they are also good models for understanding the differences in behavior on social media for topics that are highly event driven, like gun control, and those that are less focused on particular events, like climate change. Figure 1 illustrates that conversation threads related to climate change are started at a fairly consistent rate over time. By contrast, Figure 2 shows that the frequency of discussions surrounding gun control spikes sharply in response to mass shooting events. The summer of 2022 saw several deadly and highly publicized mass shootings in Buffalo, Uvalde, and Chicago that generated large amounts of discourse surrounding gun control and gun violence on social media, illustrated by the steep slope from the end of May through late July. Shortly after mass shooting events though, the level of discourse abates, indicated by a flatter slope between early August until November, when a shooting event at UVA caused another spike in conversations. Event driven issues like gun control exhibit different posting behavior and user interactions that we hypothesize leads to differences in polarization and likely would require different techniques for intervening and mitigating polarization. Developing approaches for stance labeling that are both topic independent and robust to different types of discourse online is an important step for facilitating further analysis of polarization online.
Figure 1: Gantt plot of the largest Twitter conversation threads related to climate change.
Figure 2: Gantt plot of the largest Twitter conversation threads related to gun control.
III-B User-Hashtag Stance Labeling
From the set of collected tweets for a given topic, we first construct a bipartite graph connecting users with the hashtags that each user posted. The user-hashtag bipartite graph $\mathcal{G}=(V,E)$ consists of two disjoint node sets containing the users and the hashtags posted in tweets by the user set. An edge $(u,h)\in E$ indicates that a user $u$ posted a tweet containing hashtag $h$. The weight $w(u,h)$ represents the number of times the author posted a particular hashtag across all of their authored tweets. Using the bipartite user-hashtag graph, we then develop a reciprocal label propagation-based method for automatically assigning stance labels to users and hashtags.
TABLE II: Sets of seed hashtags used for heuristic stance labeling of users in the climate change and gun control datasets.
Climate Change
Believe Disbelieve
actonclimate climatechangehoax
climatecrisis globalwarminghoax
climateaction globalcooling
climateemergency globalwarmingisahoax
climateactionnow climatehoax
Gun Control
Pro Anti
guncontrolnow shallnotbeinfringed
endgunviolence righttobeararms
gunreform gunrights
Based on a small (3-5) set of seed hashtags associated with each stance, we first propagate the stance associated with each seed hashtag across the user-hashtag bipartite graph to the set of users who posted a given labeled hashtag. The importance of a hashtag’s stance to a particular destination user is weighted by the number of time the user used each hashtag, as described in Algorithm 2. Users are then assigned a label based on the number of hashtag usages for each stance group. A user will be assigned a stance label based on the count of usages of labeled hashtags for each stance group. User labels are then propagated back to the hashtags used by each user, as described in Algorithm 3. Each hashtag is then scored based on the difference between the normalized count of usages from users of each stance group. Hashtag scores are normalized via min-max normalization and the mean and standard deviation of hashtag scores is computed. Hashtags that score one or more standard deviation above or below the mean are assigned the corresponding stance label. We then repeat alternating label propagation until the model converges, i.e. the set of labeled users is the same from the prior iteration to the next.
Algorithm 1 Bipartite stance label propagation algorithm
1:Bipartite user-hashtag graph $G$, seed hashtag set $H^{0}$
2:$H^{0}\leftarrow\{(h,s)\in{\textrm{Set of labeled seed hashtags}}\}$
3:$U^{0}\leftarrow\emptyset$
4:while $U^{(k-1)}\neq U^{k}$ do
5:     $U^{k}\leftarrow propagateTagsToUsers\left(G,\,H^{(k-1)}\right)$
6:     $H^{k}\leftarrow propagateUsersToTags\left(G,\,U^{k}\right)$
7:end while
8:return $U^{K},\,H^{K}$
Algorithm 2 Propagate hashtag labels to users
1:Bipartite user-hashtag graph $G$, labeled hashtag set $H$
2:user_counts = {}
3:for $h,s\in H$ do
4:     for $u,w\in\mathcal{N}(h)$ do
[…]
6:     end for
7:end for
[…]
10:end for
11:return $U$
Algorithm 3 Propagate user labels to hashtags
1:Bipartite user-hashtag graph $G$, labeled user set $U$
2:tag_counts = {}
3:for $u,s\in U$ do
4:     for $h\in\mathcal{N}(u)$ do
[…]
6:     end for
7:end for
[…]
9:     $h_{score}=\frac{\textrm{tag\_counts}[h][s_{2}]}{|U_{s2}|}-\frac{\textrm{tag\_counts}[h][s_{1}]}{|U_{s1}|}$
10:end for
[…]
14:     if $h_{score}\geq\mu+\sigma$ then
15:         $H\leftarrow(h,s_{1})$
16:     else if $h_{score}\leq\mu-\sigma$ then
17:         $H\leftarrow(h,s_{2})$
18:     end if
19:end for
20:return $H$
III-C GNN Stance Labeling
While the user-hashtag stance labeling method can efficiently determine the stance group of users and discover important hashtags associated with each stance group, it is limited by the fact that not all users post hashtags and that hashtags consist of a small part of the overall textual content contained in tweets. Furthermore, the method does not take into account any social interactions between users. The method has high precision but relatively low recall due to the limited breadth of users that stance labeling method, we employ the user-hashtag bipartite method to generate a soft-labeled set of users from which we train a graph neural network classifier over the attributed user-user interaction graph.
We define a user interaction as a retweet, mention, reply, or quote from the author of the tweet towards another user. We then construct a weighted, signed, and attributed user-user interaction network $G=(V,E,X)$, where $V$ is the set of users and a directed edge $(u,v)_{i}\in E$ connects the author user to the target user of their interaction as defined above. The edge weight $w(u,v)_{i}$ corresponds to the sentiment of the tweet, scored by a BERTweet-based sentiment analysis model. BERTweet is a RoBERTa model fine-tuned on English tweets and trained on the SemEval 2017 corpus. To characterized the overall interactions between users, we consolidated the multiple edges between pairs of users corresponding to individual tweets by combining each directed edge $(u,v)_{i}$ into a single edge for each user pair $(u,v)$ with the composite edge weight $w(u,v)$ computed as the mean sentiment score across all tweets between two users. Each user $u\in V$ is also associated with a feature vector $x_{u}\in X$ which corresponds to the element-wise mean of the BERTweet embedding for each tweet authored by the user. We then train a graph neural network classifier over the user interaction graph, using the set of seed users labeled by the hashtag method as a training set.
We consider three implementations from two prominent families of graph neural networks convolutional and attentional. The neighbor coefficients may be fixed weights as in convolutional GNNs or may be computed implicitly, as in attentional GNNs. Specifically, we evaluate graphSAGE[13] and Graph Isomorphism Network (GIN) as examples of convolutional GNNs, and graph attention network (GAT) as an example attentional GNN.
GraphSAGE defines a message passing function as:
$\boldsymbol{\hat{h}}_{v}^{(k)}=\sigma\left(\boldsymbol{\mathrm{W}}\sum_{u\in\mathcal{\tilde{N}}(v)}\frac{1}{|\mathcal{\tilde{N}}|}\boldsymbol{h}_{u}^{(k-1)}\right)$ (1)
where $\mathcal{\tilde{N}}(v)$ is the local neighborhood of node $v$ with added self-loop, $\boldsymbol{\mathrm{W}}$ is a trainable weight matrix, $\sigma$ is a non-linear activation function. We use the mean aggregator to aggregate the neighboring node features.
Another convolutional GNN we consider is GIN. The message passing function in GIN is defined as:
$\boldsymbol{\hat{h}}_{v}^{(k)}=\mathrm{MLP}\left((1+\epsilon)\cdot\boldsymbol{h}_{v}^{(k-1)}+\sum_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{(k-1)}\right)$ (2)
Compared with the graphSAGE operator, GIN notably replaces the single layer transformation matrix with a multi-layer perceptron (MLP)[popescu2009multilayer] and uses a simple unnormalized summation aggregator over the local neighboring nodes rather than a mean aggregator.
Thirdly, we also consider a GAT layer operator, defined as:
$\boldsymbol{\hat{h}}_{v}^{(k)}=\sigma\left(\boldsymbol{\mathrm{W}}\sum_{u\in\mathcal{\tilde{N}}(v)}a_{vu}\boldsymbol{h}_{u}^{(k-1)}\right)$ (3)
where aggregation coefficient $a_{vu}$ is computed dynamically in a node-dependent manner via edge softmax attention. In this case, the GAT function is a first-order approximation of a convolution applied over the weighted adjacency matrix, where the edge weights are computed via the attention mechanism.
III-D Heuristic Label Validation
The hashtag-based heuristic is limited in its ability to label all users, covering only ()percent of the users in the gun control dataset and ()of users in the climate change dataset. Despite coverage limitations, the method works due well to its simplicity and speed for extracting an initial set of labeled users that can serve as a training set for other language-based and network-based stance labeling methods methods. A social science expert independently annotated a set of 250 and 350 users for each dataset by analyzing the top 20 most popular tweets from each user.
Model Type Model Text Content Network Structure
Random Weighted random
Transformer BERT-base-uncased
BERT-large-uncased
RoBERTa-base
RoBERTa-large
BERTweet
GPT-4
GNN GraphSAGE
GAT
GIN
TABLE III: Models utilized for user stance labeling and indication whether the models incorporate textual based information and/or network structural information.
III-E Baselines
We compare our stance labeling method against three transformer models that leverage the language content of users’ posts. We compute an embedding for each user by computing the element-wise mean of the embedding of all of the users’ tweets from BERT, RoBERTa, and BERTweet. We then train a classifier MLP head using the labels from the hashtag heuristic to predict user stance label. We evaluate the performance of both the transformer and GNN models using 5-fold cross validation on the large set of labeled users from the reciprocal stance labeling heuristic.
We further evaluate classification performance of stance labeling methods on a small set of users manually annotated by domain experts. Lastly, we compare the performance of state-of-the-art large language models (LLMs) against the aforementioned transformer-based language models and GNNs using zero-shot classification [15, 16] with GPT4 to predict user stances for the same set of users annotated manually by human experts.
We use function calling to prompt GPT-4 to provide the sentiment of content with respect to gun control or climate change. For the gun control dataset, the function call prompt is: ”The stance of the content. Is the message pro gun control, anti gun control, or neutral¿‘ with three classification options of ”pro”, ”anti”, or ”neutral”. For the climate change dataset, ”the stance of the content. Does the message capture the user’s belief in climate change, disbelief, or is it just neutral?” with classification options of ”belief”, ”disbelief”, or ”neutral”. We then aggregate the stance labeling results for each individual tweet to determine a user’s overall stance on the issue of gun control or climate change. Users who had a majority of tweets classified by one stance were labeling accordingly and those users whose tweets were labeled as all neutral are classified as undetermined. In the case where there is an equivalent number of tweets labeled as both stances, the user is labeled as undetermined.
IV Results
We conduct two sets of evaluation for method comparison: 1) cross validation over the soft-labeled users from the hashtag-labeling heuristics; and 2) human evaluation on a set of manually annotated users. With the set of manually annotated users, we also classify each user’s stance using zero-shot classification with GPT4 to evaluate the performance of state-of-the-art large language models (LLMs) as a potential alternative to manual annotation. We report the macro averaged precision, recall, and F1 scores and we use the F1 score as the primary metric due to data imbalance. We then select the best model to classify all users in the gun control and climate change datasets to obtain their stance labels. These labels are then used to conduct a case study and comparison between polarization on these topics on Twitter.
IV-A User-Hashtag Labeling
While the user-hashtag bipartite label propagation method can quickly label a number of users with high precision. It suffers from the limitation that a large proportion of users do not use hashtags, and while it is well suited for determining hashtags that are highly associated with one particular stance group as compared with the other, hashtags that are used across both stances albeit in different contexts will remain unlabeled. Table IV As such, the method has low recall across the entire user set. Despite these limitations, the method can facilitate semi-supervised training of our socially-infused text mining labeling method that leverages the textual content of user tweets and social interactions between users.
TABLE IV: Top hashtags associated with each stance group.
Climate Change
Believe Disbelieve
climatechange globalwarming
cop26 ecocredit
climate co2
actonclimate covid
climatecrisis un
climateaction russia
ipcc maga
envrionment podestaemails8
climateemergency davos2022
sustainability ukraine
Gun Control
Pro Anti
guncontrol 2a
guncontrolnow secondamendment
gunviolence guns
nra 2ndamendment
endgunviolence breaking
uvalde scotus
gunreform shallnotbeinfringed
gunreformnow selfdefense
texas usa
gop jeroneforcongress
To determine the benefits of the reciprocal hashtag labeling heuristic, we compare the performance of the GNN stance labeling models trained on the initial set of users labeled by the seed hashtags against the performance of the GNN models for predicting user stance when trained by the set of users labeled after allowing the label propagation heuristic to converge. The initial set of labeled users consists of approximately 100,000 users for the gun control dataset and 200,000 users for the climate change dataset. The set of users labeled using reciprocal label propagation consists of approximately 400,000 users for the gun control dataset and one million users for the climate change dataset. In all, the labeling heuristic resulted in a four to five fold increase in the size of the training set of labeled users available for training the GNN stance labeling models.
For the gun control dataset, training GNN models with the expanded user set from the stance labeling heuristic results in a five to ten percent increase in precision, recall, and F1 score for all three of the graph neural network architectures. By contrast, for the climate change dataset, classification performance when training with the heuristic-labeled user set declined for all GNN architectures.
IV-B Manual Labeling
We evaluate the performance of our two-stage stance labeling procedure and the benefit of including social network connectivity for stance labeling by evaluating model predictive performance against a small set of gold standard users annotated manually by domain experts. Compared with the cross-validation performance of models evaluated on the set of heuristically, performance
IV-C GPT4 Labeling
We further examine the benefit of our two-stage stance labeling by performing zero-shot classification of user stances using GPT4 for the same sample of users annotated manually. Large language models like GPT4 have shown great promise in a broad variety of natural language tasks, and we evaluate whether GPT4 can match or exceed the performance of GNN models using just textual information. We hypothesize that by incorporating the network structural information through GNNs allows models to better encode not just a user’s utterances about a topic but also there position in the network and with whom they interact; thus allowing such models to better classify user stances on particular topics.
Contrary to the results on the cross-validation, Table LABEL:tab:gpt4-results shows that GNN models outperform tesx-based models for the Climate Change dataset, while stance prediction of users in the Gun Control dataset with GNN models performed worse than text-based models.
Enriching the user representations from text-based models with social network information improves the stance labeling performance but is not able to match the classification performance of user tweets using GPT-4
Notably, for the gun control test users, twelve users were labeled with an equivalent number of tweets from both stances by GPT-4
GPT4 Labels Manual Labels
Model Prec. Recall F1 Prec. Recall F1
Weighted random 47.36 49.99 39.45 40.76 49.92 38.11
BERT-base-uncased 0. 0. 0. 70.35 73.02 72.52
BERT-large-uncased 0. 0. 0. 74.43 75.96 75.24
RoBERTa-base 0. 0. 0. 70.28 70.62 71.85
RoBERTa-large 0. 0. 0. 71.86 71.84 71.63
BERTweet 0 0 0 77.99 80.40 77.52
GPT-4 - - - 90.02 88.45 89.14
GraphSAGE 86.83 83.30 80.36 84.64 82.93 81.76
GAT 84.96 87.52 84.91 88.00 86.95 86.18
GIN 89.64 82.49 80.78 81.87 81.52 80.48
TABLE V: Results for stance labeling classification on the Gun Control dataset ($N=350$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
GPT4 Labels Manual Labels
Model Prec. Recall F1 Prec. Recall F1
Weighted random 47.36 49.99 39.45 45.26 47.55 42.87
BERT-base-uncased 0. 0. 0. 63.32 66.41 63.75
BERT-large-uncased 0. 0. 0. 67.89 71.51 69.01
RoBERTa-base 0. 0. 0. 64.44 62.28 63.93
RoBERTa-large 0. 0. 0. 67.64 68.43 67.67
BERTweet 0 0 0 60.78 52.09 43.96
GPT-4 - - - 85.73 86.25 86.51
GraphSAGE 76.83 75.30 72.36 79.37 79.05 76.88
GAT 74.96 67.52 72.91 76.71 79.83 73.50
GIN 69.64 62.49 65.78 73.94 74.42 67.57
TABLE VI: Results for stance labeling classification on the Climate Change dataset ($N=250$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
IV-D Discussion
When researching the area of online polarization, there are serious divides in how the questions are structured and the areas of focus. This divide is also reflected in how different researchers integrate the literature and findings into future work. Oftentimes, this split can be seen in how social science researchers integrate computer science research into their studies and vice versa. Much of this is focused on the idea that both research groups are missing a key methodological piece the other offers. When researching online polarization, this methodological gap can inhibit the understanding of the black box of online polarization and radicalization. Big data research can provide scalability not feasible by traditional qualitative social science research and social science expertise provides the understanding of the subtleties of these interactions on human behavior. As has been seen in the aftermath of events like radicalized mass shooters, the online behavior of these individuals is not well understood and the need to understand how and why these processes take place online needs to be examined. The integration of subject matter expertise and AI algorithmic work is necessary for the large-scale application of big data methodology into human-level polarization research and policy work. These two methodologies, the qualitative expertise provided by subject matter experts, and the quantitative capabilities of computational researchers need to be combined so that the benefits of both are fully exploited. So the question becomes, “how do we balance the benefits of both methodologies to build the accuracy of our models?” This requires a balance of the nuance provided by subject matter experts with the speed and volume provided by computation experts. The ability to combine these two approaches can greatly benefit the understanding of online polarization.
The coding of blatant or straight-forward tweets to label the stance or polarization of the sentiment should be easily handled by AI algorithms. The difficulty comes in where the tweets are subtle, nuanced, or coded. It is in this gray space that qualitative analysis provides the necessary understanding to the complexities of human experiences, behaviors, and social phenomena. To accurately code this subset of tweets or users, the underlying meanings, and contexts have to be understood since these factors can impact how the tweets are meant to be interpreted. Quantitative research offers methodological tools like narrative analysis to analyze how culturally and historically contingent the terms, beliefs and issues narrators address in their writings are (Rosenwald, and Ochberg, 1992). Using tweets is different from traditional narrative analysis since they lack the more linear coherence and structure of “master” narratives and are instead characterized by brevity, terseness, and dialogism (Boje, 2001). Considering the utilization of Twitter [X] for social movements, political campaigning, and radicalization, these small narratives created from tweets can provide insights into the range of ongoing arguments and contentions that create collective identity online (Coupland and Brown, 2004). The benefit of this approach is that it allows for the analysis of culturally, historically, locationally nuanced and complex language which is exceptionally valuable when studying areas that are traditionally hidden or stealth online, such as alternative religions/cults, criminality, black markets, or non-traditional lifestyles/ communities (i.e. furries, incels, etc.). The wealth of this data adds critical information to understanding how people discuss their thoughts and opinions about polarizing topics like gun rights or climate change.
One area of difficulty in the process of automating AI algorithms to detect stances on topics like gun control or climate change, is that the aim is to dichotomize beliefs that exist on a spectrum. The extreme ends of these beliefs are straight-forward to categorize (i.e. individuals who do not believe climate change is real or want to ban all guns) the majority of people exist in the in-between space. So the goal then becomes to qualitatively code the tipping point as people’s stance moves between pro-and anti-gun and believe or disbelieve climate change. It is in the middle space, where coding becomes the most difficult and language nuance becomes critical to coding schemes. An example of how automating stance analysis can mislabel users because of difficulties in discerning written cues like tone, connotation, coded words, or pictures. Some of these coding errors are due to a purposeful misdirection of a tweet that is clarified by the picture. For example, user 2945287090 said “I’m pro choice…are you? https://t.co/WtYW6zohSB” but the picture had them wearing a shirt that said pro-choice but had an array of guns to “choose” from. Since the picture cannot be assessed by the algorithm, these misdirections would be potentially miscoded. A more complex example of this is a tweet by user 18643437 “A man who live streamed himself on Facebook driving around Memphis shooting strangers was caught on camera smiling after being arrested by police. https://t.co/nSYQMPric3.” The picture is a Black man in the back of a police car smiling. The coded language around this is focused on the race of the shooter as a dog-whistle and the underlying understanding by pro-gun advocates that the race of the shooter indicates that this is not what is meant by “law-abiding gun owner.” It shifts the focus away from the issues around gun violence as an overarching problem and onto scapegoating Black males as criminals who enjoy violence. There are also users who frame their arguments in neutral language or the language of the opposite side. For example, user 66811907 states “Gun control advocates do not have a monopoly on outrage or sorrow. They do not have a monopoly on ideas for addressing gun violence. They do not have a monopoly wanting to fix the problem.” The larger context of this user’s tweets balance between being pro- Second Amendment and being critical to the responses to mass shooting events, such as Uvalde. As a whole, this user falls more strongly on the side of being anti- gun control but certainly has criticism of open market gun purchasing. The neutral framing can be difficult to parse since the full narrative of the users’ tweets need to be taken into account. These different examples highlight some of the language and contextual cues and codes that need to be contextualized.
The coding errors are not only on the side of AI, however. One of the key limitations to only using human-based qualitative coding is that this type of deep understanding of stances is highly time consuming and delayed. Especially in online space, the posting moves far faster than researchers can code by hand. For mis-coded users, there are hundreds to thousands of tweets for one user. This issue means that oftentimes qualitative researchers either focus on a subset of tweets or a subset of users to be able to have insights about the online discussions of topics like guns or climate change. When examining a subset of tweets to establish stance, there will be errors. Since the coder is only looking at a section of the tweets, often randomly chosen, may not be an accurate representation of the user’s online stance. The balance between efficiency and accuracy means that there will be mis-coded information that could be “fixed” by looking at all the available data. The more narrow focus of qualitative researchers can limit the generalizability of these findings. Which for topics like polarization and radicalization, the limited generalizability then in turn limits the impact on policy and practice. The ability to integrate big-data/ AI methodology can be a force multiplier for qualitative researchers as a mixed methods approach that can improve the generalizability of these findings.
It should also be noted, as was mentioned earlier, that these stances exist on a spectrum and therefore there will be individuals who could have opposite coded stances by even two human qualitative coders. Since the researcher is making a qualitative decision of when a person’s stance moves between two ends, the middle ground can often be coded either way. For example, is a person who owns guns for hunting but pushes for very strong hand gun control measures, pro gun or anti gun? Does wanting strong background checks or other restrictions inherently make someone pro/anti gun? Or does someone who has a strong stance against their personal gun ownership but also states they do not think background checks are necessary or valuable (i.e. guns aren’t for me but I don’t care about others), pro or anti gun control? Even in the context of a large body of tweets, it may not be fully clear how this stance should be coded and is not necessarily an error on either side but rather a decision reflecting this balance. An example is user 66811907 who has a large body of tweets that different coders could code differently “This your friendly reminder that New York already bans ”assault weapons,” that the shooter procured his NY-compliant ”non-assault weapon” legally, and that the cops who stopped him showed up with their ”assault weapon” version of the same gun to combat a civilian threat. https://t.co/KLmqwQM73f” and “Rifles of any kind account for fewer annual homicides than hands and feet. Mass shooters don’t need pistol grips, etc., to cause horrific carnage, but on the flip side those features can be quite useful for lawful owners using them in self-defense. That’s why cops carry them, too.” This example is just one that highlights that neither AI nor human coding is without issues, error, or disagreement. This is seen in the reporting of inter-rater reliability in qualitative coding and could be a useful measure to add to big data analysis.
The effective coding of large data sets with an aim toward integrating these two methodologies should be the goal of research aimed at informing the larger research field and policy. Without the purposeful integration of these two methodologies, much of this research, and its impact, will continue to be limited. Both guns and climate change have a range of historical contexts, political connotations and expectations, intersections with race and gender, and sometimes conspiracy theories that can play a role in how people talk about these ideas online and inform their stances. It is in this space that the subject matter expertise can provide that contextual information that can inform the work of AI algorithms. This concept has been discussed through the idea of grounding AI with use-case specific information that is outside of the training data. The social science insights provide the ground truth necessary to properly ground AI focused on complex human stances to ensure that the AI model is acknowledging the real-world context, to minimizing errors and maximizing the relevance of the generated outputs.
IV-E Discussion
When researching the area of online polarization, there are serious divides in how the questions are structured and the areas of focus. This divide is also reflected in how different researchers integrate the literature and findings into future work. Oftentimes, this split can be seen in how social science researchers integrate computer science research into their studies and vice versa. Much of this is focused on the idea that both research groups are missing a key methodological piece the other offers. When researching online polarization, this methodological gap can inhibit the understanding of the black box of online polarization and radicalization. Big data research can provide scalability not feasible by traditional qualitative social science research and social science expertise provides the understanding of the subtleties of these interactions on human behavior. As has been seen in the aftermath of events like radicalized mass shooters, the online behavior of these individuals is not well understood and the need to understand how and why these processes take place online needs to be examined. The integration of subject matter expertise and AI algorithmic work is necessary for the large-scale application of big data methodology into human-level polarization research and policy work. These two methodologies, the qualitative expertise provided by subject matter experts, and the quantitative capabilities of computational researchers need to be combined so that the benefits of both are fully exploited. So the question becomes, “how do we balance the benefits of both methodologies to build the accuracy of our models?” This requires a balance of the nuance provided by subject matter experts with the speed and volume provided by computation experts. The ability to combine these two approaches can greatly benefit the understanding of online polarization.
The coding of blatant or straight-forward tweets to label the stance or polarization of the sentiment should be easily handled by AI algorithms. The difficulty comes in where the tweets are subtle, nuanced, or coded. It is in this gray space that qualitative analysis provides the necessary understanding to the complexities of human experiences, behaviors, and social phenomena. To accurately code this subset of tweets or users, the underlying meanings, and contexts have to be understood since these factors can impact how the tweets are meant to be interpreted. Quantitative research offers methodological tools like narrative analysis to analyze how culturally and historically contingent the terms, beliefs and issues narrators address in their writings are (Rosenwald, and Ochberg, 1992). Using tweets is different from traditional narrative analysis since they lack the more linear coherence and structure of “master” narratives and are instead characterized by brevity, terseness, and dialogism (Boje, 2001). Considering the utilization of Twitter [X] for social movements, political campaigning, and radicalization, these small narratives created from tweets can provide insights into the range of ongoing arguments and contentions that create collective identity online (Coupland and Brown, 2004). The benefit of this approach is that it allows for the analysis of culturally, historically, locationally nuanced and complex language which is exceptionally valuable when studying areas that are traditionally hidden or stealth online, such as alternative religions/cults, criminality, black markets, or non-traditional lifestyles/ communities (i.e. furries, incels, etc.). The wealth of this data adds critical information to understanding how people discuss their thoughts and opinions about polarizing topics like gun rights or climate change.
One area of difficulty in the process of automating AI algorithms to detect stances on topics like gun control or climate change, is that the aim is to dichotomize beliefs that exist on a spectrum. The extreme ends of these beliefs are straight-forward to categorize (i.e. individuals who do not believe climate change is real or want to ban all guns) the majority of people exist in the in-between space. So the goal then becomes to qualitatively code the tipping point as people’s stance moves between pro-and anti-gun and believe or disbelieve climate change. It is in the middle space, where coding becomes the most difficult and language nuance becomes critical to coding schemes. An example of how automating stance analysis can mislabel users because of difficulties in discerning written cues like tone, connotation, coded words, or pictures. Some of these coding errors are due to a purposeful misdirection of a tweet that is clarified by the picture. For example, user 2945287090 said “I’m pro choice…are you? https://t.co/WtYW6zohSB” but the picture had them wearing a shirt that said pro-choice but had an array of guns to “choose” from. Since the picture cannot be assessed by the algorithm, these misdirections would be potentially miscoded. A more complex example of this is a tweet by user 18643437 “A man who live streamed himself on Facebook driving around Memphis shooting strangers was caught on camera smiling after being arrested by police. https://t.co/nSYQMPric3.” The picture is a Black man in the back of a police car smiling. The coded language around this is focused on the race of the shooter as a dog-whistle and the underlying understanding by pro-gun advocates that the race of the shooter indicates that this is not what is meant by “law-abiding gun owner.” It shifts the focus away from the issues around gun violence as an overarching problem and onto scapegoating Black males as criminals who enjoy violence. There are also users who frame their arguments in neutral language or the language of the opposite side. For example, user 66811907 states “Gun control advocates do not have a monopoly on outrage or sorrow. They do not have a monopoly on ideas for addressing gun violence. They do not have a monopoly wanting to fix the problem.” The larger context of this user’s tweets balance between being pro- Second Amendment and being critical to the responses to mass shooting events, such as Uvalde. As a whole, this user falls more strongly on the side of being anti- gun control but certainly has criticism of open market gun purchasing. The neutral framing can be difficult to parse since the full narrative of the users’ tweets need to be taken into account. These different examples highlight some of the language and contextual cues and codes that need to be contextualized.
The coding errors are not only on the side of AI, however. One of the key limitations to only using human-based qualitative coding is that this type of deep understanding of stances is highly time consuming and delayed. Especially in online space, the posting moves far faster than researchers can code by hand. For mis-coded users, there are hundreds to thousands of tweets for one user. This issue means that oftentimes qualitative researchers either focus on a subset of tweets or a subset of users to be able to have insights about the online discussions of topics like guns or climate change. When examining a subset of tweets to establish stance, there will be errors. Since the coder is only looking at a section of the tweets, often randomly chosen, may not be an accurate representation of the user’s online stance. The balance between efficiency and accuracy means that there will be mis-coded information that could be “fixed” by looking at all the available data. The more narrow focus of qualitative researchers can limit the generalizability of these findings. Which for topics like polarization and radicalization, the limited generalizability then in turn limits the impact on policy and practice. The ability to integrate big-data/ AI methodology can be a force multiplier for qualitative researchers as a mixed methods approach that can improve the generalizability of these findings.
It should also be noted, as was mentioned earlier, that these stances exist on a spectrum and therefore there will be individuals who could have opposite coded stances by even two human qualitative coders. Since the researcher is making a qualitative decision of when a person’s stance moves between two ends, the middle ground can often be coded either way. For example, is a person who owns guns for hunting but pushes for very strong hand gun control measures, pro gun or anti gun? Does wanting strong background checks or other restrictions inherently make someone pro/anti gun? Or does someone who has a strong stance against their personal gun ownership but also states they do not think background checks are necessary or valuable (i.e. guns aren’t for me but I don’t care about others), pro or anti gun control? Even in the context of a large body of tweets, it may not be fully clear how this stance should be coded and is not necessarily an error on either side but rather a decision reflecting this balance. An example is user 66811907 who has a large body of tweets that different coders could code differently “This your friendly reminder that New York already bans ”assault weapons,” that the shooter procured his NY-compliant ”non-assault weapon” legally, and that the cops who stopped him showed up with their ”assault weapon” version of the same gun to combat a civilian threat. https://t.co/KLmqwQM73f” and “Rifles of any kind account for fewer annual homicides than hands and feet. Mass shooters don’t need pistol grips, etc., to cause horrific carnage, but on the flip side those features can be quite useful for lawful owners using them in self-defense. That’s why cops carry them, too.” This example is just one that highlights that neither AI nor human coding is without issues, error, or disagreement. This is seen in the reporting of inter-rater reliability in qualitative coding and could be a useful measure to add to big data analysis.
The effective coding of large data sets with an aim toward integrating these two methodologies should be the goal of research aimed at informing the larger research field and policy. Without the purposeful integration of these two methodologies, much of this research, and its impact, will continue to be limited. Both guns and climate change have a range of historical contexts, political connotations and expectations, intersections with race and gender, and sometimes conspiracy theories that can play a role in how people talk about these ideas online and inform their stances. It is in this space that the subject matter expertise can provide that contextual information that can inform the work of AI algorithms. This concept has been discussed through the idea of grounding AI with use-case specific information that is outside of the training data. The social science insights provide the ground truth necessary to properly ground AI focused on complex human stances to ensure that the AI model is acknowledging the real-world context, to minimizing errors and maximizing the relevance of the generated outputs.
V Conclusion
Acknowledgment
The research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-22-1-0035. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.